OpenStudy (anonymous):

The state of Wyoming has roughly the shape of a rectangle that is 3.6 x 10^2 miles long and 2.8 x 10^2 miles high. What is the approximate area of Wyoming? Hint: The area of a rectangle is product of its length and width.

OpenStudy (nottim):

what have you done so far.

OpenStudy (anonymous):

Do you multiply 3.6x10^2 and 2.8x10^2???

OpenStudy (anonymous):

\[\left( 3.6\times10^2 \right)\times \left( 2.8\times10^2 \right)=?\]

OpenStudy (anonymous):

I got 100800

OpenStudy (anonymous):

correct ,write units as well.